# On the Curvature of Regge Metrics

### v1

@@ Abstract
We use moving frame techniques to derive a notion of curvature for a class of piecewise-smooth Riemannian metrics called Regge metrics, showing that it is a measure that simultaneously satisfies the (weak) Cartan structure equations and the appropriate gauge transformation law. It turns out that this distributional curvature is equivalent to existing notions of densitized distributional curvature. We also investigate more closely the $n=2$ case, where we prove the Gauss-Bonnet theorem for Regge metrics.
@@ 1 Introduction
Since Tullio Regge’s introduction of what are today called Regge metrics—discontinuous metrics with tangential-tangential continuity—they have found widespread use in numerical models of general relativity [17, 12], continuum mechanics [15, 12, 16], and more. In all of these applications, one of the key features of a Regge metric is that a coherent notion of curvature can be defined. Regge’s original paper [17] discussed piecewise-constant metrics on simplicial meshes, where the scalar curvature is given simply by the angle defect around codimension-2 interfaces. Further investigations by Cheeger, Müller, and Schrader [3] proved that a broader class of curvatures, called the Lipschitz-Killing curvatures, converges in measure to their smooth counterparts in the piecewise-flat setting.
@@
In this paper we will focus not on the numerical convergence of distributional curvature, but rather on deriving the “correct” definition of distributional curvature from first principles. The formulas (1) and (22) are equivalent to that for the densitized distributional Riemann curvature described in [8].
@@
Importantly, a core part of the definition of a compatible frame involves blow-ups of polytopes, essentially to ensure that the frame is regular enough to permit integration by parts, even though it is discontinuous. The idea of using blow-ups to define geometric invariants of Regge metrics is not new, appearing in [1, p. 2] for much the same reason why we use them here. In Section 2, after deriving properties that such a frame must have, we use them to obtain an expression for $f^{*}\Omega_{\mathrm{dist}}$ that behaves as it should. We use blow-ups to help compute $f^{*}\Omega_{\mathrm{dist}}$ in Sections 2.2-2.3, and we also use them in Section 2.5 to derive a frame-independent expression for the distributional curvature, which is more practical for real computations, especially on manifolds that are not necessarily parallelizable.
@@
Our main results are stated in Theorems 1 and 2. Roughly speaking, they say that if an $n$-dimensional manifold is equipped with a (curvilinear) polyhedral mesh, a Regge metric $g$, and a compatible frame $f$, then the structure equations for the connection one-form and curvature two-form can be given meaning in a distributional sense. Furthermore, the distributional curvature two-form $f^{*}\Omega_{\mathrm{dist}}$, when reinterpreted as an $\mathrm{End}(TM)$-valued two-form, is a functional $\hat{R}_{\mathrm{dist}}$ that acts on any $\mathrm{End}(TM)$-valued (n2)(n-2)-form $\hat{\phi}$ with sufficient regularity via
@@
$\left\langle\left\langle\hat{R}_{\mathrm{dist}},\hat{\phi}\right\rangle\right\rangle=\sum_{T\subset M}\int_{\mathring{T}}\left\langle\hat{R}\wedge\hat{\phi}\right\rangle+\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left[\!\left[\left\langle\hat{\mathrm{I\!I}}_{e}\wedge i_{e}^{*}\hat{\phi}\right\rangle\right]\!\right]+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\left\langle\hat{\Theta}_{p}\wedge i_{p}^{*}\hat{\phi}\right\rangle$.
@@
Here, the sums over $T$, $e$, and $p$ are sums over polytopes of codimension 0, 1, and 2, respectively, and the maps $i_{e}^{*}$ and $i_{p}^{*}$ are pullbacks under the inclusions $e\hookrightarrow M$ and $p\hookrightarrow M$. The notation $\![\cdot$]\!] specifies the jump of a multi-valued quantity over the submanifold $e=T\cap T^{\prime}$, which in this case is simply the sum $\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge\hat{\phi}^{T}\right\rangle+\left\langle\hat{\mathrm{I\!I}}_{e}^{T^{\prime}}\wedge\hat{\phi}^{T^{\prime}}\right\rangle$. The quantities $\hat{R}$, $\hat{\mathrm{I\!I}}_{e}^{T}$, and $\hat{\Theta}_{p}$ are $\mathrm{End}(TM)$-valued $2$-forms, 11-forms, and 0-forms that encode the curvature tensor, second fundamental form, and angle defect, respectively. The operation $\left\langle\cdot\wedge\cdot\right\rangle$ takes a pair of $\mathrm{End}(TM)$-valued forms and wedges their form parts and applies a nondegenerate pairing to their endomorphism parts.
@@
Constructing compatible frames is nontrivial, and we provide an existence proof in Section 3. Lastly in Section 4 we investigate the specialization to two dimensions, where the Gauss curvature measure can be defined independently of the frame, and we prove that there is a suitable generalization of the Gauss-Bonnet theorem.
@@
Our results are stated for extremely general oriented parallelizable manifolds and meshes. This is partly because the presentation is actually not much more complicated for a general polyhedral mesh, since many delicate analytical conditions for Stokes’ theorem must be dealt with even in the simplest cases, and existing literature treating this subject has already been developed to a very high level of generality. One caveat is that we rely on the existence of blow-ups of polytopes, but we have only been able to locate an existence theorem for blow-ups of simplices in the literature. An explanation of the types of manifolds we use in this paper, and references to relevant literature, can be found in the appendix.
@@ 1.1 Background: Geometry in a Moving Frame
Consider an oriented, parallelizable Whitney $n$-manifold $M$ furnished with a smooth Riemannian metric $g$. The frame bundle of $M$, denoted $\mathcal{F}_{GL}(M)$, is the sub-bundle of $TM\times\dots\times TM=(TM)^{n}$ such that the fiber over each point $x\in M$ is the set of ordered bases $F=(F_{1},\dots,F_{n})$ for $T_{x}M$. The orthonormal frame bundle of $M$, denoted $\mathcal{F}_{O}(M)$, is the sub-bundle of $\mathcal{F}_{GL}(M)$ such that the fiber over each point is the set of ordered bases such that $\left\langle F_{i},F_{j}\right\rangle=\delta_{ij}$ for all $i,j$. A frame is a section of the frame bundle, and an orthonormal frame is a section of the orthonormal frame bundle.
@@
It can be shown that $\eta$ defined this way is smooth by taking a coordinate neighborhood $U\subset M$, which defines a smooth section $s:U\to\mathcal{F}_{GL}(U)$ by $s_{i}:=\frac{\partial}{\partial x_{i}}$. Every point $(x,F)\in\mathcal{F}_{GL}(X)$ is then equal to $(x,s\cdot h)$ for some matrix $h\in GL(n)$, which means $\mathcal{F}_{GL}(X)$ has smooth coordinates $(x^{i},h^{j}_{k})$. When expressed in these coordinates, $\eta$ is equal to $h^{-1}dh$. Thus $\eta$ is called the (left-invariant) Maurer-Cartan form, ubiquitous in the theory of the geometry of Lie groups and symmetric spaces ¡citations here¿.
@@
It must be noted that, in this section, we have treated $M$ as if there can exist a global frame $f:M\to\mathcal{F}_{GL}(M)$. In many cases, this is not possible. Manifolds that do support a global frame are called parallelizeable. However, every point in a manifold is contained in a neighborhood which is parallelizeable. Since curvature is a local property, this means the parallelizeable case is really the most interesting. Parallelizeability is also different from being topologically trivial; for instance, all tori are parallelizeable, as are odd-dimensional spheres.
@@ 1.2 Meshes and Regge Metrics
We will use shorthand to simplify the presentation. Again, $M$ will be assumed to be a smooth, oriented, “nice” $n$-manifold (see the appendix for precise definitions). We will also assume that $M$ is parallelizeable, with the idea in mind that more topologically complicated manifolds can be obtained by gluing together finitely many parallelizeable ones; for instance a sphere can be obtained by gluing together two disks. $M$ is equipped with a countable mesh $\{\Delta\}$ whose union is $M$, where each set $\Delta\subset M$ is the image of a closed (nd)(n-d)-dimensional convex polytope $\hat{\Delta}\subset\mathbb{R}^{n-d}$ under a smooth embedding for some $d\leq n$. If $d=0$, then the embedding must in addition be positively oriented. We will abuse terminology and call each $\Delta$ a polytope, even though it is technically the image of a polytope under a smooth embedding.
@@
The mesh must satisfy some axioms. Each face of a polytope in the mesh is also a polytope of the mesh. The intersection of two polytopes must be either a shared face of both polytopes or empty. Additionally, the mesh must respect the stratification structure of $M$. Being a “nice” manifold, the boundary $\partial M$ can be decomposed into strata $S_{d}(M)$, which consists of those points $x\in\partial M$ which are contained in a submanifold of $\partial M$ that is of codimension $d$, but not one that is of codimension d1d-1. We will require that if the relative interior of a polytope $\Delta_{d}$ intersects $S_{d^{\prime}}(M)$, then $\mathring{\Delta}_{d}\subseteq S_{d^{\prime}}(M)$. This prevents pathological tangencies at the boundary, and it means that the closure of each stratum of $M$ inherits its own mesh decomposition.
A Regge metric $g$ for the manifold $M$ and mesh $\{\Delta\}$ consists of a $C^{2}$ Riemannian metric $g_{\Delta}$ for each polytope $\Delta$, with the property that if $\Delta^{\prime}$ is a face of $\Delta$ then $i_{\Delta^{\prime}}^{*}g_{\Delta}=g_{\Delta^{\prime}}$, where $i_{\Delta^{\prime}}:\Delta^{\prime}\hookrightarrow\Delta$ is the inclusion map. In other words, the tangential-tangential components of $g$ are continuous across any shared face. Ideally, one would like to remove the positive-definiteness restriction and work with arbitrary pseudo-Riemannian geometries. We have attempted to avoid using the positive-definiteness whenever possible, but nonetheless this is still a very necessary assumption to make. In subsequent sections, $M$, $\{\Delta\}$, and $g$ are implicit.
We make frequent use of a “wedge inner product” between forms that take values in $\mathfrak{gl}(n)$. We can define a nondegenerate symmetric bilinear form $\left\langle\cdot,\cdot\right\rangle$ on the vector space $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$ by setting $\left\langle A,B\right\rangle=\mathrm{Tr}(AB)$. When restricted to $\mathfrak{so}(n)$, it is a negative-definite inner product. This means we can define a product
@@
and extending multilinearly. This is not a true inner product, but it is nondegenerate and bilinear, and it is completely independent of any metric structure, which makes it desirable for our use case. It also has some useful symmetries. One we will use often is that $\left\langle\mathrm{Ad}(h)(\alpha)\wedge\beta\right\rangle=\left\langle\alpha\wedge\mathrm{Ad}(h^{-1})(\beta)\right\rangle$ for any $h\in GL(n)$.
@@ 2 Derivation of the Distributional Riemann Curvature
The distributional Riemann curvature tensor associated to an orthonormal frame $f$, which we will denote by $f^{*}\Omega_{\mathrm{dist}}$, is a linear functional which associates a number to each smooth, compactly supported $\mathfrak{so}(n)$-valued (n2)(n-2)-form $\phi$ which vanishes when pulled back to $\partial M$. The frame $f$ is made up of $C^{2}$ frames on the interior of each codimension-0 polytope $T$, so $f=\bigsqcup_{T\subset M}f^{T}$ and each $f^{T}:\mathring{T}\to\mathcal{F}_{O}(\mathring{T})$ is a $C^{2}$ section. We will also take as a definition that $f^{*}\omega$ and $f^{*}[\omega,\omega]$ are the piecewise-$C^{2}$ forms defined as $f^{*}\omega|_{\mathring{T}}:={f^{T}}^{*}\omega$ and $f^{*}[\omega,\omega]|_{\mathring{T}}:={f^{T}}^{*}[\omega,\omega]$, where $\omega\in C^{1}\Omega^{1}(\mathcal{F}_{O}(T);\mathfrak{so}(n))$ is the usual connection one-form. The distributional exterior derivative of $f^{*}\omega$ is the linear functional defined by
@@
What we aim to do is find conditions on the frame $f$ such that the right-hand side of (5) can be efficiently computed, does not depend on derivatives of $\phi$, and transforms like a tensor.

### v2

@@ Abstract
We use moving frame techniques to derive a notion of curvature for a class of piecewise-smooth Riemannian metrics called Regge metrics, showing that it is a measure that simultaneously satisfies the (weak) Cartan structure equations and the appropriate gauge transformation law. It turns out that this distributional curvature is equivalent to existing notions of densitized distributional curvature. We also investigate more closely the two-dimensional case, where we prove the Gauss-Bonnet theorem for Regge metrics.
@@ 1 Introduction
Since Tullio Regge’s introduction of what are today called Regge metrics—discontinuous metrics with tangential-tangential continuity—they have found widespread use in numerical models of general relativity [22, 17], continuum mechanics [20, 17, 21], and more. In many of these applications, one of the key features of a Regge metric is that curvature measures can be defined which converge in measure to their smooth counterparts. Regge’s original paper [22] discussed piecewise-constant metrics on simplicial meshes, where the scalar curvature is given simply by the angle defect around codimension-2 interfaces. Further investigations by Cheeger, Müller, and Schrader [5] proved that a broader class of curvatures, called the Lipschitz-Killing curvatures, converges in measure to their smooth counterparts in the piecewise-flat setting.
@@
In this paper we will focus not on the numerical convergence of distributional curvature, but rather on deriving the “correct” definition of distributional curvature from first principles. The formulas (4) and (22) are equivalent to that for the densitized distributional Riemann curvature described in [13]. We are not the first to pursue a derivation of the correct definition from first principles, especially for the scalar curvature in the piecewise-flat setting; see [7, 6, 3, 4] and the references therein for other perspectives on this.
@@
Importantly, a core part of the definition of a compatible frame involves blow-ups of polytopes, essentially to ensure that the frame is regular enough to permit integration by parts, even though it is discontinuous. The idea of using blow-ups to define geometric invariants of Regge metrics is not new, appearing in [2, p. 2] for much the same reason why we use it here. In Section 2, after deriving properties that such a frame must have, we use them to obtain an expression for $f^{*}\Omega_{\mathrm{dist}}$ that behaves as it should. We use blow-ups to help compute $f^{*}\Omega_{\mathrm{dist}}$ in Sections 2.2-2.3, and we also use them in Section 2.5 to derive a frame-independent expression for the distributional curvature, which is more practical for real computations, especially on manifolds that are not necessarily parallelizable.
@@
Our main results are stated in Theorems 4 and 9. Roughly speaking, they say that if an $n$-dimensional manifold $M$ is equipped with a (curvilinear) polyhedral mesh, a Regge metric $g$, and a compatible frame $f$, then the structure equations for the connection one-form and curvature two-form can be given meaning in a distributional sense. Furthermore, the distributional curvature two-form $f^{*}\Omega_{\mathrm{dist}}$, when reinterpreted as an $\mathrm{End}(TM)$-valued two-form, is a functional $\hat{R}_{\mathrm{dist}}$ that acts on any $\mathrm{End}(TM)$-valued (n2)(n-2)-form $\hat{\phi}$ with sufficient regularity via
@@
$\left\langle\left\langle\hat{R}_{\mathrm{dist}},\hat{\phi}\right\rangle\right\rangle=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle\hat{R}\wedge\hat{\phi}\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left[\!\left[\left\langle\hat{\mathrm{I\!I}}_{e}\wedge i_{\mathring{e}}^{*}\hat{\phi}\right\rangle\right]\!\right]+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\left\langle\hat{\Theta}_{p}\wedge i_{\mathring{p}}^{*}\hat{\phi}\right\rangle$.
@@
Here, the sums over $T$, $e$, and $p$ are sums over polytopes of codimension 0, 1, and 2, respectively, and the maps $i_{\mathring{e}}^{*}$ and $i_{\mathring{p}}^{*}$ are pullbacks under the inclusions $\mathring{e}\hookrightarrow M$ and $\mathring{p}\hookrightarrow M$. The notation $\![\cdot$]\!] specifies the jump of a multi-valued quantity over the submanifold $e=T\cap T^{\prime}$, which in this case is simply the difference $\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\hat{\phi}|_{T}\right\rangle-\left\langle\hat{\mathrm{I\!I}}_{e}^{T^{\prime}}\wedge i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}\right\rangle$. The quantities $\hat{R}$, $\hat{\mathrm{I\!I}}_{e}^{T}$, and $\hat{\Theta}_{p}$ are $\mathrm{End}(TM)$-valued $2$-forms, 11-forms, and 0-forms that encode the curvature tensor, second fundamental form, and angle defect, respectively. The operation $\left\langle\cdot\wedge\cdot\right\rangle$ takes a pair of $\mathrm{End}(TM)$-valued forms and wedges their form parts and applies a nondegenerate pairing to their endomorphism parts.
@@
Constructing compatible frames is nontrivial, and we provide an existence proof in Section 3. Lastly, in Section 4 we investigate the specialization to two dimensions, where the Gauss curvature measure can be defined independently of the frame, and we prove that there is a suitable generalization of the Gauss-Bonnet theorem.
@@
Our results are stated for extremely general oriented parallelizable manifolds and meshes. This is partly because the presentation is actually not much more complicated for a general polyhedral mesh, since many delicate analytical conditions for Stokes’ theorem must be dealt with even in the simplest cases, and existing literature treating this subject has already been developed to a very high level of generality. One caveat is that we rely on the existence of blow-ups of polytopes which are also polytopes, but we have only been able to locate an existence theorem for blow-ups of simplices in the literature. An explanation of the types of manifolds we use in this paper, and references to relevant literature, can be found in the appendix.
@@ 1.1 Background: Geometry in a Moving Frame
Consider an oriented, parallelizable polyhedral $n$-manifold $M$ furnished with a smooth Riemannian metric $g$. The frame bundle of $M$, denoted $\mathcal{F}_{GL}(M)$, is the sub-bundle of $TM\times\dots\times TM=(TM)^{n}$ such that the fiber over each point $x\in M$ is the set of ordered bases $F=(F_{1},\dots,F_{n})$ for $T_{x}M$. The orthonormal frame bundle of $M$, denoted $\mathcal{F}_{O}(M)$, is the sub-bundle of $\mathcal{F}_{GL}(M)$ such that the fiber over each point is the set of ordered bases such that $\left\langle F_{i},F_{j}\right\rangle=\delta_{ij}$ for all $i,j$. A frame is a section of the frame bundle, and an orthonormal frame is a section of the orthonormal frame bundle.
@@
It can be shown that $\eta$ defined this way is smooth by taking a coordinate neighborhood $U\subset M$, which defines a smooth section $s:U\to\mathcal{F}_{GL}(U)$ by $s_{i}:=\frac{\partial}{\partial x_{i}}$. Every point $(x,F)\in\mathcal{F}_{GL}(U)$ is then equal to $(x,s\cdot h)$ for some matrix $h\in GL(n)$, which means $\mathcal{F}_{GL}(U)$ has smooth coordinates $(x^{i},h^{j}_{k})$. When expressed in these coordinates, $\eta$ is equal to $h^{-1}dh$. Thus $\eta$ is called the (left-invariant) Maurer-Cartan form, ubiquitous in the theory of the geometry of Lie groups and symmetric spaces [23].
@@
It must be noted that, in this section, we have treated $M$ as if there can exist a global frame $f:M\to\mathcal{F}_{GL}(M)$. In many cases, this is not possible. Manifolds that do support a global frame are called parallelizable. However, every point in a manifold is contained in a neighborhood which is parallelizable. Since curvature is a local property, this means the parallelizable case is really the most interesting. Parallelizability is also different from being topologically trivial; for instance, spheres of dimension 0, 11, 33, and $7$ and products thereof are all parallelizable.
@@ 1.2 Meshes and Regge Metrics
$M$ will be assumed to be a smooth, oriented, polyhedral $n$-manifold (see the appendix for precise definitions). We will also assume that $M$ is parallelizable, with the idea in mind that more topologically complicated manifolds can be obtained by gluing together finitely many parallelizable ones; for instance a sphere can be obtained by gluing together two disks. $M$ is equipped with a countable mesh $\{\Delta\}$ whose union is $M$, where each set $\Delta\subset M$ is the image of a closed (nd)(n-d)-dimensional convex polytope $\hat{\Delta}\subset\mathbb{R}^{n-d}$ under a smooth embedding for some $d\leq n$. If $d=0$, then the embedding must in addition be positively oriented. We will abuse terminology and call each $\Delta$ a polytope, even though it is technically the image of a polytope under a smooth embedding.
@@
The mesh must satisfy some axioms. Each face of a polytope in the mesh is also a polytope of the mesh. The intersection of two polytopes must be either a shared face of both polytopes or empty. Additionally, the mesh must respect the stratification structure of $M$. Being a polyhedral manifold, the boundary $\partial M$ can be decomposed into strata $S_{d}(M)$, which consists of those points $x\in\partial M$ which are contained in a submanifold of $\partial M$ that is of codimension $d$, but not one that is of codimension d1d-1. We will require that if the relative interior of a polytope $\Delta_{d}$ intersects $S_{d^{\prime}}(M)$, then $\mathring{\Delta}_{d}\subseteq S_{d^{\prime}}(M)$. This prevents pathological tangencies at the boundary, and it means that the closure of each stratum of $M$ inherits its own mesh decomposition.
A Regge metric $g$ for the manifold $M$ and mesh $\{\Delta\}$ consists of a $C^{2}$ Riemannian metric $g^{\Delta}$ for each polytope $\Delta$, with the property that if $\Delta^{\prime}$ is a face of $\Delta$ then $i_{\Delta^{\prime}}^{*}g^{\Delta}=g^{\Delta^{\prime}}$, where $i_{\Delta^{\prime}}:\Delta^{\prime}\hookrightarrow\Delta$ is the inclusion map. In other words, the tangential-tangential components of $g$ are continuous across any shared face. Ideally, one would like to remove the positive-definiteness restriction and work with arbitrary pseudo-Riemannian geometries. We have attempted to avoid using the positive-definiteness whenever possible, but nonetheless this is still a very necessary assumption to make. In subsequent sections, $M$, $\{\Delta\}$, and $g$ are implicit.
We make frequent use of a “wedge inner product” between forms that take values in $\mathfrak{gl}(n)$. We can define a nondegenerate symmetric bilinear form $\left\langle\cdot,\cdot\right\rangle$ on the vector space $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$ by setting $\left\langle A,B\right\rangle:=\mathrm{Tr}(AB)$. When restricted to $\mathfrak{so}(n)$, it is a negative-definite inner product. This means we can define a product
@@
and extending multilinearly. This is not a true inner product, but it is nondegenerate and bilinear, and it is completely independent of any metric structure, which makes it desirable for our use case. It also has some useful symmetries. One we will use often is that $\left\langle(\mathrm{Ad}(h)(A)\otimes\alpha)\wedge(B\otimes\beta)\right\rangle=\left\langle(A\otimes\alpha)\wedge(\mathrm{Ad}(h^{-1})(B)\otimes\beta)\right\rangle$ for any $h\in GL(n)$. When applying the adjoint action (or any other map $\mathfrak{gl}(n)\to\mathfrak{gl}(n)$) to a Lie algebra part of a Lie algebra valued form, we will abuse notation slightly by applying it to the whole form.
@@ 2 Derivation of the Distributional Riemann Curvature
The distributional Riemann curvature tensor associated to an orthonormal frame $f$, which we will denote by $f^{*}\Omega_{\mathrm{dist}}$, is a linear functional which associates a number to each smooth, compactly supported $\mathfrak{so}(n)$-valued (n2)(n-2)-form $\phi$ which vanishes when pulled back to $\partial M$. The frame $f$ is made up of $C^{2}$ frames on the interior of each codimension-0 polytope $T$, so $f=\bigsqcup_{T\subseteq M}f^{T}$ and each $f^{T}:\mathring{T}\to\mathcal{F}_{O}(\mathring{T})$ is a $C^{2}$ section. We will also take as a definition that $f^{*}\omega$ and $f^{*}[\omega,\omega]$ are the piecewise-$C^{2}$ forms defined as $f^{*}\omega|_{\mathring{T}}:={f^{T}}^{*}\omega$ and $f^{*}[\omega,\omega]|_{\mathring{T}}:={f^{T}}^{*}[\omega,\omega]$, where $\omega\in C^{1}\Omega^{1}(\mathcal{F}_{O}(\mathring{T});\mathfrak{so}(n))$ is the usual connection one-form. The distributional exterior derivative of $f^{*}\omega$ is the linear functional defined by
@@
What we aim to do is find conditions on the frame $f$ such that the right-hand side of (5) can be efficiently computed, is bounded by a multiple of the supremum norm of $\phi$, and transforms like a tensor.